OpenStudy (anonymous):

Find all solutions of the equation 2sin2x−cosx=1 in the interval [0,2π) Find all solutions of the equation 2sin2x−cosx=1 in the interval [0,2π) @Mathematics

OpenStudy (anonymous):

hey it is 2sin^2(x)-cox=1 in the interval [0,2pi)

OpenStudy (anonymous):

Good - might have been a bit more difficult. Since sin^2x+cos^2x=1 we have sin^2x=1-cos^2x. Using this in our equation gives: 2(1-cos^2x)-cosx=1. Multiplying out the LHS and subtracting 1 over gives: 2-2cos^2x-cosx-1=0 Or, rearranging terms: -2cos^2x-cosx+1=0. Factoring gives: (-2cosx+1)(cosx+1)=0. This gives two equations: -2cosx+1=0 => cosx=1/2 => x=pi/4 and 7pi/4 in [0, 2pi) and cosx+1=0 => cosx=-1 .=> x=pi in [0, 2pi). Solutions: x=pi/4, pi, and 7pi/4 in [0, 2pi).

OpenStudy (anonymous):

what about this one Find all solutions of the equation sec^(2)x−2=0. The answer is A+Bkπ where k is any integer and 0<A<π/2

OpenStudy (anonymous):

find A and B